Stryker Industries received an offer from an exporter for 15,000 units of product at $17.50 per unit. The acceptance of the offer will not affect normal production or domestic sales prices. The following data is available: Domestic unit sales price $20 Unit manufacturing costs: Variable 11 Fixed 1 What is the amount of income or loss from acceptance of the offer?

  • $97,500 income
  • B. $94,500 loss
  • C. $37,500 income
  • D. $37,500 loss
Answer #1

Revenue from special order

= 15,000 * 17.50 = 262,500

Relevant cost of making that order

= 15,000 * 11 = 165,000

Income (Loss) = 262,500 - 165,000

= 97,500 income